OpenStudy (anonymous):

From 5 employees at a company, a group of 3 employees will be chosen to work on a project. How many different groups of 3 employees can be chosen?

OpenStudy (anonymous):

\(\dbinom{5}{3}\) or \(_5C_3\) read "five choose 3" and computed via \[\binom{5}{3}=\frac{5\times 4}{2}\]

OpenStudy (anonymous):

I did n!/r!(n!-r!) and got the wrong answer. I might be doing it wrong.

OpenStudy (anonymous):

don't get married to the formula

OpenStudy (anonymous):

first of all, it should be clear that 5 choose 3 is the same as 5 choose 2 is that clear or no?

OpenStudy (anonymous):

oh, i see, your formula is wrong, but no matter you should use it anyway

OpenStudy (anonymous):

*SHOULDN'T

OpenStudy (anonymous):

I'm sorry I haven't studied math for over a year and trying to refresh my math skills for placement testing hha

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

because if you think, rather than use a formula, this is really easy

OpenStudy (anonymous):

ok:D

OpenStudy (anonymous):

you have five people (or whatever) and you want to select 2 of them make a fraction, you have two slots to fill up top as \(5\times 4\) and in the denominator you put \(2\times 1\) which really only need to write \(2\)

OpenStudy (anonymous):

so you have \(_5C_2=\frac{5\times 4}{2}=5\times 2=10\)

OpenStudy (anonymous):

wanna do another one that is a bit harder?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets compute \(\binom{10}{7}\)

OpenStudy (anonymous):

the first thing to realize is that choosing 7 out of 10 is identical to choosing 3 out of 10 is that clear or no?

OpenStudy (anonymous):

yes that's clear.

OpenStudy (anonymous):

ok so instead we will compute \[\binom{10}{3}\] now we have 3 slots to fill in the numerator starting at 10 to give a numerator of \(10\times 9\times 8\) and a denominator starting at 3 giving \(3\times 2\)

OpenStudy (anonymous):

so \[\binom{10}{7}=\binom{10}{3}=\frac{10\times 9\times 8}{3\times 2}\] cancel first multiply last

OpenStudy (anonymous):

you get \[\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

ohh it's getting back to me now, this is paired with sigma as well right?

OpenStudy (anonymous):

I remember doing something like this back in high school but i just totally forgot the name of it

OpenStudy (anonymous):

if you want to write a formula, it is \[\binom{n}{k}=\frac{n!}{k!(n-k)!}\] but the formula is silly to use because it just tells you all you need to cancel

OpenStudy (anonymous):

there are formula that involve sigmas, which is about addition, and some are related to combinations, but it is not really the same

OpenStudy (anonymous):

ok, now I understand how the answer was 10. Thank you so much! I don't want to place in a prerequisite class at my college because I already took pre-calc back in high school..plus the classes that come before pre-calc in college don't even count towards graduation.

OpenStudy (anonymous):

hahah, so thank you very much!!

OpenStudy (anonymous):

yeah it is easy to get stuck in those annoying low math courses, and then they make them harder than they need to good luck yw